# Reflections to set-theoretic solutions of the Yang-Baxter equation 11footnotemark: 1

### v2

@@ Introduction
Our study is motivated by a recent description of left non-degenerate solutions in terms of left shelves structures provided in [11, Theorem 2.15]. Indeed, any left non-degenerate solution can be written in terms of a family of automorphisms of the left shelf associated with it, i.e., the Drinfel’d twist, and of the shelf operation itself. This naturally suggests splitting the study of reflections into their behaviour with respect to the maps $\lambda_{x}\in\operatorname{Aut}\left(X,\triangleright_{r}\right)$ and left multiplications $L_{x}$. Hence, reflections further motivate the investigation of the endomorphism monoid of shelves.
We first focus on reflections for solutions that are only left or right non-degenerate since, in our analysis, they turn out to be different from each other. Following some existing approaches mentioned before, we pay particular attention to $\lambda$-centralizing and $\rho$-invariant reflections and characterize them in the left and right non-degenerate cases in 2.1 and in 3.1, respectively. Besides, we reformulate the results achieved in terms of the respective multiplication groups of their associated shelf structure (see 2.10 and 3.4). Then, we deal with the non-degenerate bijective case, by combining the observations made separately for the left and right cases, and developing a theory for such solutions. In this context, in Appendix A, we gather some numerical data on reflections for bijective non-degenerate solutions associated with skew braces, dividing them by classes of belonging.
The last section is devoted to constructing reflections starting from known ones. More specifically, considering construction techniques for arbitrary solutions, as the matched product [2] and the strong semilattice of solutions [3], we define reflections for the new solution obtained, starting from reflections on each solution involved in the construction itself. Moreover, we focus on the twisted union, a method for obtaining involutive non-degenerate solutions provided in [15]. In particular, we first extend this techinque to arbitrary solutions and then we show how to construct a reflection for new solutions determined.
@@ 1.1 Basics on the reflection equation
[28, Example 1.4] Let $X$ be a set, $f,g\in\operatorname{Sym}_{X}$ such that $fg=gf$, and consider the bijective non-degenerate solution $r(x,y)=(f(y),g(x))$ belonging to the class of Lyubashenko’s solutions [13]. A map $\kappa:X\to X$ is a reflection for $(X,r)$ if and only if $\kappa$ commutes with $fg$. In particular, if $(X,r)$ involutive, i.e., $g=f^{-1}$ , any map $\kappa:X\to X$ is a reflection for $(X,r)$.
@@
In [28, Lemma 1.7], it is introduced the following notation to study the reflections of a solution $(X,r)$.
@@
In [26, Theorem 3.2], it is shown that for involutive non-degenerate solutions, it is sufficient to check only one of the relations (T) and (Q).
@@
The authors in [28, Theorem 1.9] give sufficient conditions for a map $\kappa$ to be a reflection for an involutive left non-degenerate solution. To recall such a result, let us introduce the following notion.
@@
We underline that, in general, not all the reflections for an involutive left non-degenerate solution are $\lambda$-centralizing (see, for instance, $\kappa_{2}$ in 1.5). Regarding the right non-degenerate case, we recall the following result contained in [26, Theorem 3.2].
@@ 1.2 Some basics on shelf structures
A shelf homomorphism between two shelves is defined as any map preserving the shelf operations on each underlying set. Hereafter, we will denote with $\operatorname{End}(X,\triangleright)$ and $\operatorname{Aut}(X,\triangleright)$ the endomorphism monoid and the automorphism group of a given shelf $(X,\triangleright)$, respectively. If we denote with $L_{x}:X\to X,y\mapsto x\triangleright y$ the operator of left multiplication by $x$, for all $x\in X$, then equation (1) is equivalent to require that $L_{x}$ is a shelf homomorphism.
@@
for all $x,y\in X$, then the structure $(X,\triangleright_{r})$ is a left shelf, called the left shelf associated to the solution $(X,r)$. The solution associated with $\left(X,\triangleright_{r}\right)$ is called the left-derived solution of $(X,r)$.
@@
According to [11, Theorem 2.15], it is possible to describe any left (or right) non-degenerate solution in terms of its associated shelf. We recall such a result in the direction useful for our purposes.
@@ 2 Reflections for left non-degenerate solutions
It directly follows from 2.1 and (1). ∎
The property 2. of 2.1 further simplifies if we assume that $L_{x}$ is bijective, for all $x\in X$, i.e., $(X,\triangleright_{r})$ is a rack. According to [11, Corollary 2.16], this is the case of a left non-degenerate solution which is also bijective.
@@
$\forall\,x\in X\quad\kappa L_{x}=\kappa L_{\kappa\left(x\right)}\,.$
In particular, $\kappa$ is an involutive reflection for $(X,r)$ if and only if $\kappa$ is an involutive automorphism of $\left(X,\triangleright_{r}\right)$.
@@
Let $\kappa_{1},\kappa_{2}$ be $\lambda$-centralizing reflections of a bijective left non-degenerate solution $(X,r)$. If $\kappa_{1}\kappa_{2}=\kappa_{2}\kappa_{1}$, then $\kappa_{1}\kappa_{2}$ is a reflection of $(X,r)$. The converse is not true since in 1.5 it holds that $\kappa_{2}\kappa_{3}=\kappa_{2}$ and $\kappa_{3}\kappa_{2}=\kappa_{4}$. Moreover, in general, the commutation assumption can not be relaxed since for a given solution there can exist $\lambda$-centralizing reflections $\kappa$ and $\omega$ whose composition is not again a reflection (see 1.3-3.).
@@
The 2.7 allows us to easily find examples of reflections of the solutions associated with racks.
@@
The proof of the fact that $r_{\kappa}$ is a bijective and non-degenerate solution is a consequence of [11, Theorem 2.15] and 2.7. Moreover, the rest of the claim follows from 2.5. ∎
Notice that the solution $(X,r_{\kappa})$ in 2.9 is equivalent to $(X,r_{\triangleright})$ if and only if $\kappa=\operatorname{id}_{X}$. An easy calculation shows that $\triangleright=\triangleright_{r_{\kappa}}$ and therefore $r_{\triangleright}$ and $r_{\kappa}$ share the same left rack structure. According to [11, Lemma 2.25], it follows that $r_{\triangleright}$ and $r_{\kappa}$ are Drinfel’d-isomorphic, with the explicit Drinfel’d-isomorphism $\Phi:X\times X\to X\times X$ given by $\Phi(x,y)=(x,\kappa(y))$, for all $x,y\in X$.
@@
Let $X=\mathbb{Z}_{n}$ and $(X,\triangleright)$ be the dihedral quandle of order $n\in\mathbb{N}$. In [14, Theorem 2.1] the authors prove that $\operatorname{Aut}(X)=\operatorname{Aff}(\mathbb{Z}_{n})$, the group of affine transformations of $\mathbb{Z}_{n}$. With an analogous proof one can compute $\operatorname{End}(X)$ and show that
@@
In light of 2.7, a map $\kappa:X\to X$ is a reflection for $(X,r_{\triangleright})$ if and only if there exist $a,b\in\mathbb{Z}_{n}$ such that $\kappa=f_{b,a}$ and it satisfies $\kappa L_{x}=\kappa L_{\kappa(x)}$, for all $x\in\mathbb{Z}_{n}$. Therefore, a straightforward computation shows that $\kappa\in\mathcal{K}(X,r_{\triangleright})$ if and only if
@@
An analogous procedure can be carried on to determine that the only reflections for the left derived solution associated to the infinite dihedral quandle are the identity map and the maps of arbitrary constant value.
Let us conclude this section with some considerations on idempotent left non-degenerate solutions.
@@
    $\kappa:X\to X$ is a $\lambda$-centralizing reflection if and only if  $\kappa^{2}=\kappa$;
@@
Let us initially observe that $\rho_{y}\left(x\right)=\lambda^{-1}_{\lambda_{x}\left(y\right)}\lambda_{x}\left(y\right)$, for all $x,y\in X$. Hence, the first statement easily follows from Lemma 1.6. For the second one, note that (Q) is satisfied if and only if $\kappa\rho_{y}=\rho_{y}$, for every $y\in X$. Consequently, (T) holds if and only if $\lambda_{x}\kappa=\lambda_{x}$, for every $x\in X$, and, since $r$ is left non-degenerate, we get $\kappa=\operatorname{id}_{X}$. ∎
Note that the previous result is similar to that in the bijective case, i.e., 2.5, since the shelf associated with any idempotent solution is that in 1.15-$2$ with $f=\operatorname{id}_{X}$, namely, the trivial shelf.
@@ 3 Reflections for right-non degenerate solutions
Such a map $\kappa$ has to satisfy $\kappa(2y)=2y-\kappa(0)$ and $\kappa(2y-1)=2y-\kappa(1)$, for all $y\in X$, therefore it is parametrized by its values in 00 and $1$. To determine their values we notice that $02\kappa(0)=$ 0 and $02(\kappa(1)-1)=$ 0, as a consequence of (2). Therefore, all reflections for $(X,r_{\triangleleft})$ can be classified as follows:
@@
As a direct consequence of 3.6, if $(X,\triangleleft)$ is a right rack then $Z(\operatorname{RMlt}(X,\triangleleft))\subseteq\mathcal{K}(X)$. In particular, if the right multiplication group is abelian - as in the case of a cyclic rack - then every right multiplication provides a reflection for the associated solution.
@@ 4 Reflections for bijective non-degenerate solution
Let us start with some considerations on involutive non-degenerate solutions $(X,r)$. One can easily show that any $\rho$-invariant reflection can be described in terms of the retract solution of $(X,r)$. Recall that the equivalence relation of retraction on $X$ is defined by setting $x\sim y\iff\lambda_{x}=\lambda_{y}$, for all $x,y\in X$. Thus, there is a natural induced solution $\operatorname{Ret}(X,r)$ on the quotient $\tilde{X}:=X/\sim$, called the retract of $(X,r)$, which itself is involutive and non-degenerate. For more information, we refer the reader to the standard literature (see [15, Section 3]).
@@
The retraction relation for an involutive non-degenerate solution $(X,r)$ can be equivalently defined by requiring that $x\sim y\iff\rho_{x}=\rho_{y}$, for all $x,y\in X$. This directly follows by the identity $\rho_{x}=T\lambda^{-1}_{x}T^{-1}$, for all $x\in X$, where $T:X\to X$ is the bijective map defined by $T(x)=\rho_{x}^{-1}(x)$, for all $x\in X$ (cf. [15, Proposition 2.2]).
@@
Let $(X,r)$ be an involutive non-degenerate solution. Then, they hold:
@@
The two statements are a direct consequence of 1.10. ∎
Remark 4.3.
Let $\left(X,r\right)$ be an involutive non-degenerate solution with $X$ finite and assume that $\kappa:X\to X$ is a $\rho$-invariant map. If $r$ is irretractable, then $\kappa\in\mathcal{K}_{\text{inj}}(X,r)$. Indeed, by 1.10, $\kappa$ is a reflection. Moreover, if $x,y\in X$ are such that $\kappa\left(x\right)=\kappa\left(y\right)$, then $\rho_{x}=\rho_{\kappa\left(x\right)}=\rho_{\kappa\left(y\right)}=\rho_{y}$, hence, by irretractability, $x=y$. As a consequence, since $X$ is finite, then $\kappa\in\mathcal{K}_{\text{bij}}(X,r)$.
Now, we aim to combine the results obtained in the previous sections and analyze what happens in the special case of bijective non-degenerate solutions. To do so, in [25, cf. Proposition 3.9], it is shown the existing link between left and right structure racks.
@@
If we specialize the 4.6 to a solution $(X,r)$ of derived type, it does not yield a novel result, as we explain below. One can analogously reason if $(X,r)$ is a solution of right derived type.
@@
If $(X,r)$ is a solution whose left derived solution is associated with a left rack $(X,\triangleright)$, then its right derived solution $(X,r_{\triangleleft})$ has the form $r_{\triangleleft}(x,y)=(y\triangleleft_{r}x,x)=(x\triangleright y,x)$, for all $x,y\in X$, whence $\triangleleft_{r}=\triangleright^{\text{op}}$. Therefore, the map $T$ from Lemma 4.4 actually is an automorphism of $(X,\triangleright)$. If we notice that $\operatorname{Aut}(X,r)=\operatorname{Aut}(X,\triangleright)$ then 4.6 simply recovers the observation made in 1.2.
@@
First assume that $\kappa\in\mathcal{K}(X,r)$. Then, since by assumption $\kappa$ is $\rho$-invariant, 4.6-2 ensures that $\kappa$ commutes with all $R_{x}$. As a consequence, for all $x,y\in X$ the following holds:
@@
proving the first part. Conversely, the assertion easily follows from 4.6-2. since the hypothesis yields that $\kappa$ satisfies $\kappa R_{x}=R_{x}\kappa$, for every $x\in X$. ∎
In a general setting and inspired by [26, Proposition 3.5], we show how to obtain reflections for arbitrary solutions starting from a reflection and maps that are $\lambda,\rho$-centralizing and $\lambda,\rho$-invariant.
@@
Notice that if $(X,r)$ is an involutive solution, then, for the maps $\varphi,\psi$, the assumptions of being $\lambda$-invariant and $\rho$-centralizing are redundant. Hence, the result in [26, Proposition 3.5] is a special case of 4.9.
@@
For the first three examples we employed 4.6 to verify that they actually are reflections. Moreover, there are 64 which are neither $\lambda$-centralizing nor $\rho$-invariant. However, by applying 4.9, we can recover each of them in the form $\varphi\kappa\psi$ for some $\lambda,\rho$-centralizing and $\lambda,\rho$-invariant maps $\varphi$, $\psi$ (different from the identity) and for some $\kappa\in\mathcal{K}(X,r)$. For instance, if $\kappa_{4}$ is as in the example above, then $\kappa_{4}=\varphi\kappa\psi$ where $\varphi=21436587=\psi$ and $\kappa=12556611$.
Finally, we provide a class of bijective non-degenerate solutions, including the solution in 2.9, whose reflections are not yet necessarily $\lambda$-centralizing or $\rho$-invariant.
@@
Let $(X,\triangleright)$ a left rack and $\varphi\in\operatorname{Aut}(X,\triangleright)$. Then, [11, Theorem 2.15] ensures that the map $r(x,y)=\left(\varphi(y),L_{y}\varphi^{-1}(x)\right)\,$, for all $x,y\in X$, is a bijective non-degenerate solution. An easy application of Lemma 1.6 allows to determine that

### v3

@@ Introduction
Our study is motivated by a recent description of left non-degenerate solutions in terms of left shelves structures provided in [12, Theorem 2.15]. Indeed, any left non-degenerate solution can be written in terms of a family of automorphisms of the left shelf associated with it, i.e., the Drinfel’d twist, and of the shelf operation itself. This naturally suggests splitting the study of reflections into their behaviour with respect to the maps $\lambda_{x}\in\operatorname{Aut}\left(X,\triangleright_{r}\right)$ and left multiplications $L_{x}$. Hence, reflections further motivate the investigation of the endomorphism monoid of shelves.
We first focus on reflections for solutions that are only left or right non-degenerate since, in our analysis, they turn out to be different from each other. Following some existing approaches mentioned before, we pay particular attention to $\lambda$-centralizing and $\rho$-invariant reflections and characterize them in the left and right non-degenerate cases in 2.1 and in 3.1, respectively. Besides, we reformulate the results achieved in terms of the respective multiplication groups of their associated shelf structure (see 2.11 and 3.4). Then, we deal with the non-degenerate bijective case, by combining the observations made separately for the left and right cases, and developing a theory for such solutions. In this context, in Appendix A, we gather some numerical data on reflections for bijective non-degenerate solutions associated with skew braces, dividing them by classes of belonging.
The last section is devoted to constructing reflections starting from known ones. More specifically, considering construction techniques for arbitrary solutions, as the matched product [2] and the strong semilattice of solutions [3], we define reflections for the new solution obtained, starting from reflections on each solution involved in the construction itself. Moreover, we focus on the twisted union, a method for obtaining involutive non-degenerate solutions provided in [16]. In particular, we first extend this technique to arbitrary solutions and then we show how to construct a reflection for new solutions determined.
@@ 1.1 Basics on the reflection equation
[29, Example 1.4] Let $X$ be a set, $f,g\in\operatorname{Sym}_{X}$ such that $fg=gf$, and consider the bijective non-degenerate solution $r(x,y)=(f(y),g(x))$ belonging to the class of Lyubashenko’s solutions [14]. A map $\kappa:X\to X$ is a reflection for $(X,r)$ if and only if $\kappa$ commutes with $fg$. In particular, if $(X,r)$ involutive, i.e., $g=f^{-1}$, any map $\kappa:X\to X$ is a reflection for $(X,r)$.
@@
In [29, Lemma 1.7], it is introduced the following notation to study the reflections of a solution $(X,r)$.
@@
In [27, Theorem 3.2], it is shown that for involutive non-degenerate solutions, it is sufficient to check only one of the relations (T) and (Q).
@@
The authors in [29, Theorem 1.9] give sufficient conditions for a map $\kappa$ to be a reflection for an involutive left non-degenerate solution. To recall such a result, let us introduce the following notion.
@@
We underline that, in general, not all the reflections for an involutive left non-degenerate solution are $\lambda$-centralizing (see, for instance, $\kappa_{2}$ in 1.5). Regarding the right non-degenerate case, we recall the following result contained in [27, Theorem 3.2].
@@ 1.2 Some basics on shelf structures
A shelf homomorphism between two shelves is defined as any map preserving the shelf operations on each underlying set. Hereafter, we will denote with $\operatorname{End}(X,\triangleright)$ and $\operatorname{Aut}(X,\triangleright)$ the endomorphism monoid and the automorphism group of a given shelf $(X,\triangleright)$, respectively. If we denote with $L_{x}:X\to X,y\mapsto x\triangleright y$ the operator of left multiplication by $x$, for all $x\in X$, then equation (1) is equivalent to requiring that $L_{x}$ is a shelf endomorphism.
@@
for all $x,y\in X$, then the structure $(X,\triangleright_{r})$ is a left shelf, called the left shelf associated to the solution $(X,r)$. The solution associated with $\left(X,\triangleright_{r}\right)$ is called the left derived solution of $(X,r)$.
@@
According to [12, Theorem 2.15], it is possible to describe any left (or right) non-degenerate solution in terms of its associated shelf. We recall such a result in the direction useful for our purposes.
@@ 2 Reflections for left non-degenerate solutions
It directly follows from 2.1 and the self-distributivity property (1). ∎
The property 2. of 2.1 further simplifies if we assume that $L_{x}$ is bijective, for all $x\in X$, i.e., $(X,\,\triangleright_{r})$ is a rack. According to [12, Corollary 2.16], this is the case of a left non-degenerate solution which is also bijective.
@@
$\forall\,x\in X\quad\kappa L_{x}=\kappa L_{\kappa\left(x\right)}\,.$ (2)
In particular, the following hold:
  1. 1.
    if $\kappa\in\mathcal{K}(X,r)$ is involutive, then $\kappa$ is an involutive automorphism of $\left(X,\,\triangleright_{r}\right)$.
  2. 2.
    $\kappa\in\mathcal{K}(X,r)$ is idempotent, i.e., $\kappa^{2}=\kappa$, if and only if $\kappa$ is an idempotent of $\operatorname{End}\left(X,\,\triangleright_{r}\right)$;
Proof.
The result is an easy consequence of 2.1. Moreover, $1$ is obvious. Finally, to prove $2$, it is sufficient to observe that if we assume that $\kappa:X\to X$ is an idempotent endomorphism of $(X,\,\triangleright_{r})$, then the following holds, for all $x\in X$:
$\kappa L_{\kappa(x)}=L_{\kappa^{2}(x)}\kappa=L_{\kappa(x)}\kappa=\kappa L_{x}\,,$
thus finishing the proof. ∎
Observe that, in general, the converse of 2.5-$1$ is not true (see 2.15).
@@
Reflection maps share an unexpected connection with averaging operators on racks. According to [7], if $(X,\,\triangleright)$ is a rack then a map $A:X\to X$ is a (left) averaging operator on $(X,\,\triangleright)$ if  $A\left(x\right)\triangleright A\left(y\right)=A\left(A\left(x\right)\triangleright y\right)$, for all $x,y\in X$. Evidently, if $A\in\operatorname{End}(X,\,\triangleright)$, then $A$ is averaging operator on  $(X,\,\triangleright)$ if and only if $A$ satisfies the identity (2) in 2.5.
Remark 2.7.
Let $\kappa_{1},\kappa_{2}$ be $\lambda$-centralizing reflections of a bijective left non-degenerate solution $(X,r)$. If $\kappa_{1}\kappa_{2}=\kappa_{2}\kappa_{1}$, then $\kappa_{1}\kappa_{2}$ is a reflection of $(X,r)$. The converse is not true since in 1.5 we have that $\kappa_{2}\kappa_{3}=\kappa_{2}$ and $\kappa_{3}\kappa_{2}=\kappa_{4}$. Moreover, in general, the commutation assumption can not be relaxed since for a given solution there can exist $\lambda$-centralizing reflections $\kappa$ and $\omega$ whose composition is not again a reflection (see 1.3-3.).
@@
The 2.8 allows us to easily find examples of reflections of the solutions associated with racks.
@@
The proof of the fact that $r_{\kappa}$ is a bijective and non-degenerate solution is a consequence of [12, Theorem 2.15] and 2.8. Moreover, the rest of the claim follows from 2.5. ∎
Notice that the solution $(X,r_{\kappa})$ in 2.10 is equivalent to $(X,r_{\triangleright})$ if and only if $\kappa=\operatorname{id}_{X}$. An easy calculation shows that $\triangleright=\triangleright_{r_{\kappa}}$ and therefore $r_{\triangleright}$ and $r_{\kappa}$ share the same left rack structure. According to [12, Lemma 2.25], it follows that $r_{\triangleright}$ and $r_{\kappa}$ are Drinfel’d-isomorphic, with the explicit Drinfel’d-isomorphism $\Phi:X\times X\to X\times X$ given by $\Phi(x,y)=(x,\kappa(y))$, for all $x,y\in X$.
@@
Let $X=\mathbb{Z}_{n}$ and $(X,\triangleright)$ be the dihedral quandle of order $n\in\mathbb{N}$. In [15, Theorem 2.1] the authors prove that $\operatorname{Aut}(X)=\operatorname{Aff}(\mathbb{Z}_{n})$, the group of affine transformations of $\mathbb{Z}_{n}$. With an analogous proof one can compute $\operatorname{End}(X)$ and show that
@@
In light of 2.8, a map $\kappa:X\to X$ is a reflection for $(X,r_{\triangleright})$ if and only if there exist $a,b\in\mathbb{Z}_{n}$ such that $\kappa=f_{b,a}$ and it satisfies $\kappa L_{x}=\kappa L_{\kappa(x)}$, for all $x\in\mathbb{Z}_{n}$. Therefore, a straightforward computation shows that $\kappa\in\mathcal{K}(X,r_{\triangleright})$ if and only if
@@
An analogous procedure can be carried out to determine that the only reflections for the left derived solution associated to the infinite dihedral quandle are the identity map and the maps of arbitrary constant value.
Remark 2.15.
Note that the converse of 2.5-$1$ is not true. Indeed, the map
$f_{1,-1}:\mathbb{Z}_{3}\to\mathbb{Z}_{3}\,,f_{1,-1}(x)=1-x$
is an involutive automorphism of the dihedral quandle on $\mathbb{Z}_{3}$ (see 2.14), but it is not a reflection for the associated left derived solution since, in this case, the unique reflection is the identity map.
Let us conclude this section with some considerations on idempotent left non-degenerate solutions. Observe that the shelf associated with any idempotent solution $(X,r)$ is the trivial shelf given by $L_{x}=x$, for every $x\in X$.
@@
    a $\lambda$-centralizing map $\kappa:X\to X$ is a reflection if and only if  $\kappa^{2}=\kappa$;
@@
Let us initially observe that the first statement is a direct consequence of 2.1. For the second one, note that (Q) is satisfied if and only if $\kappa\rho_{y}=\rho_{y}$, for every $y\in X$. Consequently, (T) holds if and only if $\lambda_{x}\kappa=\lambda_{x}$, for every $x\in X$, and, since $r$ is left non-degenerate, we get $\kappa=\operatorname{id}_{X}$. ∎
@@ 3 Reflections for right-non degenerate solutions
Such a map $\kappa$ has to satisfy $\kappa(2y)=2y-\kappa(0)$ and $\kappa(2y-1)=2y-\kappa(1)$, for all $y\in X$, therefore it is parametrized by its values in 00 and $1$. To determine their values, we notice that $02\kappa(0)=$ 0 and $02(\kappa(1)-1)=$ 0, as a consequence of (3). Therefore, all reflections for $(X,r_{\triangleleft})$ can be classified as follows:
@@
As a direct consequence of 3.6, if $(X,\triangleleft)$ is a right rack, then $Z(\operatorname{RMlt}(X,\triangleleft))\subseteq\mathcal{K}(X)$. Moreover, if $R_{a}$ is a reflection then $R_{a}\in\operatorname{End}\left(X,\triangleleft\right)$. In particular, if the right multiplication group is abelian - as in the case of a cyclic rack - then every right multiplication provides a reflection for the associated solution.
@@ 4 Reflections for bijective non-degenerate solution
Let us start with some considerations on involutive non-degenerate solutions $(X,r)$. One can easily show that any $\rho$-invariant reflection can be described in terms of the retract solution of $(X,r)$. Recall that the equivalence relation of retraction on $X$ is defined by setting $x\sim y\iff\lambda_{x}=\lambda_{y}$, for all $x,y\in X$. Thus, there is a natural induced solution $\operatorname{Ret}(X,r)$ on the quotient $\tilde{X}:=X/\sim$, called the retract of $(X,r)$, which itself is involutive and non-degenerate. In the sequel, $[x]_{\sim}$ denotes the equivalence class of any $x\in X$. For more information, we refer the reader to the standard literature (see [16, Section 3]).
@@
The retraction relation for an involutive non-degenerate solution $(X,r)$ can be equivalently defined by requiring that $x\sim y\iff\rho_{x}=\rho_{y}$, for all $x,y\in X$. This follows directly by the identity $\rho_{x}=T\lambda^{-1}_{x}T^{-1}$, for all $x\in X$, where $T:X\to X$ is the bijective map defined by $T(x)=\rho_{x}^{-1}(x)$, for all $x\in X$ (cf. [16, Proposition 2.2]).
@@
Let $(X,r)$ be an involutive non-degenerate solution. Then, the following hold:
@@
The two statements are a direct consequence of 1.10. We added the proof in the file. ∎
Now, we aim to combine the results obtained in the previous sections and analyze what happens in the special case of bijective non-degenerate solutions. To do so, in [26, cf. Proposition 3.9], it is shown the existing link between left and right structure racks (see also [26, Remark 3.10]).
@@
If we specialize the 4.5 to a solution $(X,r)$ of derived type, it does not yield a novel result, as we explain below. One can argue analogously if $(X,r)$ is a solution of right derived type.
@@
If $(X,r)$ is a solution whose left derived solution is associated with a left rack $(X,\triangleright)$, then its right derived solution $(X,r_{\triangleleft})$ has the form $r_{\triangleleft}(x,y)=(y\triangleleft_{r}x,x)=(x\triangleright y,x)$, for all $x,y\in X$, whence $\triangleleft_{r}=\triangleright^{\text{op}}$. Therefore, the map $T$ from Lemma 4.3 actually is an automorphism of $(X,\triangleright)$. If we notice that $\operatorname{Aut}(X,r)=\operatorname{Aut}(X,\triangleright)$ then 4.5 simply recovers the observation made in 1.2.
@@
First, assume that $\kappa\in\mathcal{K}(X,r)$. Then, since by assumption $\kappa$ is $\rho$-invariant, 4.5-2 ensures that $\kappa$ commutes with all $R_{x}$. As a consequence, for all $x,y\in X$ the following holds:
@@
proving the first part. Conversely, the assertion easily follows from 4.5-2. since the hypothesis yields that $\kappa$ satisfies $\kappa R_{x}=R_{x}\kappa$, for every $x\in X$. ∎
In a general setting and inspired by [27, Proposition 3.5], we show how to obtain reflections for arbitrary solutions starting from a reflection and maps that are $\lambda,\rho$-centralizing and $\lambda,\rho$-invariant.
@@
Notice that if $(X,r)$ is an involutive solution, then, for the maps $\varphi,\psi$, the assumptions of being $\lambda$-invariant and $\rho$-centralizing are redundant. Hence, the result in [27, Proposition 3.5] is a special case of 4.8.
@@
For the first three examples we employed 4.5 to verify that they actually are reflections. Moreover, there are 64 which are neither $\lambda$-centralizing nor $\rho$-invariant. However, by applying 4.8, we can recover each of them in the form $\varphi\kappa\psi$ for some $\lambda,\rho$-centralizing and $\lambda,\rho$-invariant maps $\varphi$, $\psi$ (different from the identity) and for some $\kappa\in\mathcal{K}(X,r)$. For instance, if $\kappa_{4}$ is as in the example above, then $\kappa_{4}=\varphi\kappa\psi$ where $\varphi=21436587=\psi$ and $\kappa=12556611$.
Finally, we provide a class of bijective non-degenerate solutions, including the solution in 2.10, whose reflections are not necessarily $\lambda$-centralizing or $\rho$-invariant.
@@
Let $(X,\triangleright)$ a left rack and $\varphi\in\operatorname{Aut}(X,\triangleright)$. Then, [12, Theorem 2.15] ensures that the map $r(x,y)=\left(\varphi(y),L_{y}\varphi^{-1}(x)\right)\,$, for all $x,y\in X$, is a bijective non-degenerate solution. An easy application of Lemma 1.6 allows to determine that